# A Retrieval Augmented Spatio-Temporal Framework for Traffic Prediction

### v1

@@ IV-B Performance Evaluation (RQ1)
IV-C Ablation Study (RQ2)
Method MAE RMSE MAPE (%)
@@ IV-C Ablation Study (RQ2)
TABLE IV: Ablation study results for different components on the PEMS07 dataset. We compare the full RAST with 7 variants. $\downarrow$ Deg% denotes the degradation percentage.
We conducted comprehensive ablation studies to evaluate the contribution of each component within our RAST, with results presented in Table IV. The results demonstrate the significance of each component, with performance degradations across different metrics when components are removed or modified. The query generator emerges as the most critical component, with its removal causing the most substantial performance degradation of 25.6% in MAE, 15.4% in RMSE, and 42.9% in MAPE. This dramatic decline validates our design of context-aware query generation as fundamental to the framework’s effectiveness. Both spatial and temporal encoders prove indispensable, with their removal resulting in 17.2% and 21.2% MAE degradation respectively, confirming that dual-stream encoding effectively captures essential spatio-temporal dependencies.
@@
IV-D Parameter Sensitivity Analysis (RQ3)
@@ IV-E Efficiency Analysis (RQ4)
TABLE V: Memory and efficiency comparisons on large-scale datasets. Mem: CUDA memory (GB) used, Train: training time (seconds per epoch), Val: inference time (in seconds).
RAST demonstrates exceptional computational efficiency, as shown in Table V. We selected several models that demonstrated superior performance in previous experiments and evaluated them using maximum possible batch sizes (32 for SD and 16 for GBA) on a single GPU, recording three critical metrics: memory usage, training time per epoch, and inference time. While graph/attention-based models face a quadratic growing computational cost w.r.t. the number of nodes, RAST achieves the fastest training and inference speeds across both GBA and SD datasets, with training times of 154.08 and 45.53 seconds per epoch respectively, and inference times of 43.52 and 10.15 seconds. Though STGCN exhibits lower memory consumption, it requires significantly more time for both training and inference, highlighting that our model achieves better overall efficiency. Complex models like D2STGNN face severe scalability challenges on large datasets, while RAST maintains consistent performance scaling through the retrieval-augmented mechanism to capture low-predictability patterns under limited contextual capacity constraints, instead of relying on complex architectures. Through the sustainable storage and maintenance of the retrieval store, RAST maintains high performance while keeping the computational cost similar to STGCN, which verifies the effectiveness of our method under limited contextual capacity.
V Conclusion
In this paper, we introduce RAST, a universal spatio-temporal forecasting framework that leverages retrieval-augmented mechanisms to address the challenges of modeling fine-grained spatio-temporal dependencies under limited context capacity. Extensive experiments on 6 real-world traffic datasets, including traffic prediction, ablation studies, hyperparameter sensitivity analysis, and the efficiency study, demonstrate that competitive performance of RAST on the spatio-temporal forecasting task while maintaining computational efficiency. These results underscore the potential of retrieval-augmented approaches to advance spatio-temporal modeling for large-scale and heterogeneous scenarios. Our future work will focus on (i) extending RAST to broader domains including climate modeling and electricity demand forecasting, and (ii) optimizing inference efficiency across diverse pre-trained spatio-temporal graph neural networks.
@@ V Conclusion
1) Non-Spatial Methods: ARIMA [5], a classical linear autoregressive moving average model for univariate time-series forecasting; VAR [6], a vector autoregression approach capturing linear dependencies across multiple time series; SVR [54], a support vector regression method leveraging kernel tricks for nonlinear time-series prediction; LSTM [10], a recurrent neural network architecture designed for learning long-term dependencies; TCN [55], a temporal convolutional network utilizing dilated causal convolutions for sequence modeling; Transformer [56], a self-attention-based architecture allowing parallel and long-range sequence modeling.
2) Spatial-temporal GNN Methods: DCRNN [14], integrating diffusion convolution with recurrent units for effective spatio-temporal modeling on graphs; STGCN [12], combining graph convolutions and temporal convolutions for spatio-temporal feature extraction; ASTGCN [34], introducing spatial and temporal attention mechanisms for dynamic feature weighting; GWNet [15], leveraging adaptive graph structures for flexible spatial dependency learning; LSGCN [57], a lightweight spatio-temporal GCN tailored for efficiency; STSGCN [53], utilizing spatio-temporal subgraph convolutions to capture local dependencies; STFGNN [58], fusing spatial and temporal features through a unified graph neural network; STGODE [59], employing neural ordinary differential equations to model continuous spatio-temporal dynamics; DSTAGNN [28], applying dynamic spatial-temporal attention for complex dependency modeling; AGCRN [13], using adaptive graph convolution and node-specific embeddings for personalized forecasting.
3) Other Enhanced Approaches: Z-GCNETs [64], a zero-shot generalizable graph convolutional network for spatio-temporal prediction; EnhanceNet [64], an advanced spatio-temporal framework that aggregates multi-scale features for robust prediction; TAMP [61], leveraging temporal and spatial multi-head attention for adaptive representation learning; STKD [62], introducing knowledge distillation to transfer knowledge in spatio-temporal models; RPMixer [63], utilizing random projection layers within all-MLP mixer blocks to capture spatial-temporal dependencies.
@@ Appendix B Datasets Descriptions
The traffic network datasets we have used in experiments and their descriptions are as follows:
@@
The Performance Measurement System (PeMS) datasets [65] are widely used benchmarks for traffic forecasting, maintained by the California Department of Transportation. These datasets collect real-time traffic data from sensors installed across California’s freeway system, providing multi-dimensional features including traffic flow, speed, and occupancy measurements with 5-minute temporal resolution. The adjacency matrices for these datasets are constructed based on road network distances between sensor locations. In our experiments, we follow the standard protocol of using a 70%/10%/20% split for training, validation, and testing respectively, with the temporal order preserved to ensure realistic evaluation scenarios.
@@
The LargeST benchmark [25] represents a significant advancement in large-scale traffic forecasting datasets. This benchmark extends beyond traditional small-scale datasets by providing comprehensive large-scale traffic networks that better reflect real-world deployment scenarios. The LargeST datasets span multiple years (2017-2021) with 5-minute temporal resolution and include comprehensive metadata for spatial topology construction. For our evaluation, we adopt the data processing pipeline provided by the LargeST benchmark with a split ratio of 6:2:2, utilizing the 2019 data slice for consistent comparison with baseline methods.
@@ C-B Training Optimization
TABLE VI: Default Training Parameters for the Experiments.
All models are implemented in PyTorch and trained on Nvidia RTX A6000 GPUs with 48GB of memory. Table VI presents the comprehensive training parameters used for RAST across all experiments. We use the Adam optimizer with a multi-step learning rate scheduler to ensure stable convergence.
C-C Model Architecture Parameters
@@
TABLE VII: Model Architecture Parameters for RAST.
Table VII details the model architecture parameters for RAST. We leverage the Spatio-Temporal Retrieval Store based on the FAISS library for efficient similarity search in the retrieval store. The retrieval store utilizes GPU acceleration for index operations to maintain computational efficiency.
@@
The computational complexity of our retrieval mechanism is dominated by the similarity search operation, which leverages FAISS with inverted file index (IVF) to achieve $O(k\log M+kd)$ complexity, where $k$ is the number of retrieved patterns, $M$ is the memory size, and $d$ is the embedding dimension. This represents a significant improvement over the $O(N^{2})$ complexity of attention mechanisms in large graphs. The momentum-based memory update follows exponential moving average dynamics $\mathcal{M}^{(t+1)}=(1-\alpha)\mathcal{M}^{(t)}+\alpha\mathcal{E}^{(t)}$, where $\alpha$ is the update rate. This formulation ensures that the memory maintains recent patterns while preserving historically significant ones, with convergence guaranteed under standard assumptions of bounded update rates and Lipschitz continuity.
@@
Our decoupled spatial and temporal encoders employ different architectural designs optimized for their respective modalities. The temporal encoder utilizes 1D convolutions with dilation to capture multi-scale temporal patterns, while the spatial encoder employs graph convolution operations adapted to the road network topology. The multi-head attention mechanism for fusing query and retrieval embeddings uses scaled dot-product attention with temperature scaling $\tau=0.1$ to control attention distribution sharpness, with 16 attention heads chosen to balance representational capacity and computational efficiency. The memory management strategy implements a hybrid approach that combines temporal decay for patterns older than 50 epochs, similarity pruning for low-quality patterns below 0.3 similarity threshold, and capacity management that bounds memory size to 1000 patterns per bank to ensure computational efficiency while maintaining pattern diversity.
Appendix E Limitation and Future Work

### v2

@@ 4.2 Performance Evaluation (RQ1)
4.3 Ablation Study (RQ2)
Methods MAE RMSE MAPE (%)
@@ 4.3 Ablation Study (RQ2)
Table 3: Ablation study results for different components on the PEMS07 dataset. We compare the full RAST with 7 variants. $\downarrow$ Deg% denotes the degradation percentage.
We conducted comprehensive ablation studies to evaluate the contribution of each component within our RAST, with results presented in Table 3. The results demonstrate the significance of each component, with performance degradations across different metrics when components are removed or modified. The query generator emerges as the most critical component, with its removal causing the most substantial performance degradation of 25.6% in MAE, 15.4% in RMSE, and 42.9% in MAPE. This dramatic decline validates our design of context-aware query generation as fundamental to the framework’s effectiveness. Both spatial and temporal encoders prove indispensable, with their removal resulting in 17.2% and 21.2% MAE degradation respectively, confirming that dual-stream encoding effectively captures essential spatio-temporal dependencies.
@@
4.4 Parameter Sensitivity Analysis (RQ3)
@@ 4.5 Efficiency Analysis (RQ4)
Table 4: Memory and efficiency comparisons on large-scale datasets. Mem: CUDA memory (GB) used, Train: training time (seconds per epoch), Val: inference time (in seconds).
RAST demonstrates exceptional computational efficiency, as shown in Table 4. We selected several models that demonstrated superior performance in previous experiments and evaluated them using maximum possible batch sizes (32 for SD and 16 for GBA) on a single GPU, recording three critical metrics: memory usage, training time per epoch, and inference time. While graph/attention-based models face a quadratic growing computational cost w.r.t. the number of nodes, RAST achieves the fastest training and inference speeds across both GBA and SD datasets, with training times of 154.08 and 45.53 seconds per epoch respectively, and inference times of 43.52 and 10.15 seconds. Though STGCN exhibits lower memory consumption, it requires significantly more time for both training and inference, highlighting that our model achieves better overall efficiency. Complex models like D2STGNN face severe scalability challenges on large datasets, while RAST maintains consistent performance scaling through the retrieval-augmented mechanism to capture low-predictability patterns under limited contextual capacity constraints, instead of relying on complex architectures. Through the sustainable storage and maintenance of the retrieval store, RAST maintains high performance while keeping the computational cost similar to STGCN, which verifies the effectiveness of our method under limited contextual capacity.
5 Conclusion
In this paper, we introduce RAST, a universal spatio-temporal forecasting framework that integrates retrieval-augmented mechanisms to tackle fine-grained spatio-temporal dependencies under limited context capacity. Comprehensive experiments across six real-world traffic datasets, including predictive benchmarks, ablation analyses, hyperparameter studies, and efficiency evaluations. The results demonstrate that RAST delivers state-of-the-art performance while retaining favorable computational efficiency. These findings highlight the promise of retrieval-augmented designs as a lightweight yet powerful complement to conventional spatio-temporal architectures, especially in large-scale and heterogeneous urban scenarios. In the future, we will continue to (i) extend RAST to broader domains such as climate modeling and electricity demand forecasting, and (ii) further optimize inference efficiency and online adaptability when interfacing with diverse pre-trained spatio-temporal graph neural networks.
Acknowledgments
This work is mainly supported by the National Natural Science Foundation of China (No. 62402414). This work is also supported by the Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515011994), Guangzhou Municipal Science and Technology Project (No. 2023A03J0011), the Guangzhou Industrial Information and Intelligent Key Laboratory Project (No. 2024A03J0628), and a grant from State Key Laboratory of Resources and Environmental Information System, and Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No. 2023B1212010007).
@@ Acknowledgments
1) Non-Spatial Methods: ARIMA (box2015time), a classical linear autoregressive moving average model for univariate time-series forecasting; VAR (lutkepohl2005new), a vector autoregression approach capturing linear dependencies across multiple time series; SVR (awad2015support), a support vector regression method leveraging kernel tricks for nonlinear time-series prediction; LSTM (hochreiter1997long), a recurrent neural network architecture for learning long-term dependencies; TCN (lea2017temporal), a temporal convolutional network utilizing dilated causal convolutions for sequence modeling; Transformer (vaswani2017attention), a self-attention-based architecture allowing parallel and long-range sequence modeling; NHiTS (challu2022nhitsneuralhierarchicalinterpolation), leveraging hierarchical interpolation and multi-rate data sampling to decompose time series into trend and seasonal components; iTransformer(liu2023itransformer), applying attention over variate tokens and feed-forward networks across time to capture multivariate correlations; TimeMixer (wang2024timemixer), leveraging decomposable multiscale mixing to disentangle intricate temporal variations by aggregating seasonal information from fine-to-coarse scales and trend components from coarse-to-fine scales.
2) Spatial-temporal GNN Methods: DCRNN (li2017diffusion), integrating diffusion convolution with recurrent units for effective spatio-temporal modeling on graphs; STGCN (yu2017spatio), combining graph convolutions and temporal convolutions for spatio-temporal feature extraction; ASTGCN (guo2019attention), introducing spatial and temporal attention mechanisms for dynamic feature weighting; GWNet (wu2019graph), leveraging adaptive graph structures for flexible spatial dependency learning; LSGCN (huang2020lsgcn), a lightweight spatio-temporal GCN tailored for efficiency; STSGCN (song2020spatial), utilizing spatio-temporal subgraph convolutions to capture local dependencies; STFGNN (li2021spatial), fusing spatial and temporal features through a unified graph neural network; STGODE (fang2021spatial), employing neural ordinary differential equations to model continuous spatio-temporal dynamics; DSTAGNN (lan2022dstagnn), applying dynamic spatial-temporal attention for complex dependency modeling; AGCRN (bai2020adaptive), using adaptive graph convolution and node-specific embeddings for personalized forecasting.
3) Other Enhanced Approaches: Z-GCNETs (chen2021z), a zero-shot generalizable graph convolutional network for spatio-temporal prediction; EnhanceNet (cirstea2021enhancenet), an advanced spatio-temporal framework that aggregates multi-scale features for robust prediction; TAMP (chen2022tamp), leveraging temporal and spatial multi-head attention for adaptive representation learning; RPMixer (yeh2024rpmixer), utilizing random projection layers within all-MLP mixer blocks to capture spatial-temporal dependencies. STKD (wang2024spatial), introducing knowledge distillation to transfer knowledge in spatio-temporal models; STDN (cao2025spatiotemporal), constructing dynamic graph structures with spatio-temporal embeddings to disentangle traffic flow into trend-cyclical and seasonal components.
@@ Appendix B Datasets Descriptions
Datasets #Points #Samples #TimeSlices Timespan
PEMS03 358 9.38M 26,208 09/01/2018-11/30/2018
PEMS04 307 5.22M 16,992 01/01/2018-02/28/2018
PEMS07 883 24.92M 28,224 05/01/2017-08/31/2017
PEMS08 170 3.04M 17,856 07/01/2016-08/31/2016
SD 716 25M 35040 01/01/2019-12/31/2019
GBA 2352 82M 35040 01/01/2019-12/31/2019
Table 5: The Statistics Details of the Dataset.
The detailed statistics of traffic network datasets we have used in experiments are shown in Table 5, and their descriptions are as follows.
@@
The Performance Measurement System (PeMS) datasets (chen2001freeway) are widely used benchmarks for traffic forecasting, maintained by the California Department of Transportation. These datasets collect real-time traffic data from sensors installed across California’s freeway system, providing multi-dimensional features including traffic flow, speed, and occupancy measurements with 5-minute temporal resolution. The adjacency matrices for these datasets are constructed based on road network distances between sensor locations. In our experiments, we follow the standard protocol of using a 70%/10%/20% split for training, validation, and testing respectively, with the temporal order preserved to ensure realistic evaluation scenarios.
@@
The LargeST benchmark (liu2023largest) represents a significant advancement in large-scale traffic forecasting datasets. This benchmark extends beyond traditional small-scale datasets by providing comprehensive large-scale traffic networks that better reflect real-world deployment scenarios. The LargeST datasets span multiple years (2017-2021) with 5-minute temporal resolution and include comprehensive metadata for spatial topology construction. For our evaluation, we adopt the data processing pipeline provided by the LargeST benchmark with a split ratio of 6:2:2, utilizing the 2019 data slice for consistent comparison with baseline methods.
@@ C.2 Training Optimization
Table 6: Default Training Parameters for the Experiments.
All models are implemented in PyTorch and trained on Nvidia RTX A6000 GPUs with 48GB of memory. Table 6 presents the comprehensive training parameters used for RAST across all experiments. We use the Adam optimizer with a multi-step learning rate scheduler to ensure stable convergence.
C.3 Model Architecture Parameters
@@
Table 7: Model Architecture Parameters for RAST.
Table 7 details the model architecture parameters for RAST. We leverage the Spatio-Temporal Retrieval Store based on the FAISS library for efficient similarity search in the retrieval store. The retrieval store utilizes GPU acceleration for index operations to maintain computational efficiency.
@@
The computational complexity of our retrieval mechanism is dominated by the similarity search operation, which leverages FAISS with an inverted file index (IVF) to achieve $O(k\log M+kd)$ complexity, where $k$ is the number of retrieved patterns, $M$ is the memory size, and $d$ is the embedding dimension. This represents a significant improvement over the $O(N^{2})$ complexity of attention mechanisms in large graphs. The momentum-based memory update follows exponential moving average dynamics $\mathcal{M}^{(t+1)}=(1-\alpha)\mathcal{M}^{(t)}+\alpha\mathcal{E}^{(t)}$, where $\alpha$ is the update rate. This formulation ensures that the memory maintains recent patterns while preserving historically significant ones, with convergence guaranteed under standard assumptions of bounded update rates and Lipschitz continuity.
@@
Our decoupled spatial and temporal encoders employ different architectural designs optimized for their respective modalities. The temporal encoder utilizes 1D convolutions with dilation to capture multi-scale temporal patterns, while the spatial encoder employs graph convolution operations adapted to the road network topology. The multi-head attention mechanism for fusing query and retrieval embeddings uses scaled dot-product attention with temperature scaling $\tau=0.1$ to control attention distribution sharpness, with 4 attention heads chosen to balance representational capacity and computational efficiency. The memory management strategy implements a hybrid approach that combines temporal decay for patterns older than 50 epochs, similarity pruning for low-quality patterns below 0.3 similarity threshold, and capacity management that bounds memory size to 1000 patterns per bank to ensure computational efficiency while maintaining pattern diversity.
Appendix E Motivation of Feature Disentanglement
Spatio-temporal forecasting tasks involve two heterogeneous sources of dynamics: temporal dependencies that often exhibit multi-scale periodicity, and spatial correlations that capture localized and topology-driven interactions. In traditional architectures, these two aspects are encoded jointly in high-dimensional embeddings, which can result in entangled representations that obscure the distinct regularities within each dimension. Such entanglement not only increases the difficulty of learning fine-grained patterns but also inflates the cost of storing and retrieving contextual information. To address this limitation, we introduce a disentanglement strategy that encodes temporal and spatial features in separate but complementary streams before fusion. This allows temporal encoders to focus on cyclical variations such as rush-hour periodicity, while spatial encoders concentrate on local network connectivity. The separation reduces interference between modalities and preserves domain-specific inductive biases, enhancing both interpretability and efficiency.
Feature disentanglement also provides a foundation for the retrieval-augmented mechanism in RAST. By decoupling temporal and spatial embeddings, our framework maintains two compact retrieval stores, each specialized for one dimension. This dual-store design enables efficient vector indexing, storage, and retrieval that scales linearly with embedding size, avoiding the cubic growth associated with storing full spatio-temporal tensors. Furthermore, the disentanglement improves retrieval quality, as queries can be matched against dimension-specific historical patterns rather than against mixed high-dimensional vectors. The subsequent cross-attention fusion ensures that the complementary contributions of both dimensions are integrated, preserving the overall capacity to model coupled spatio-temporal dependencies.
Appendix F Justification for Dual-Dimensional Encoding and Retrieval
A spatio-temporal embedding $H\in\mathbb{R}^{N\times T\times d}$ typically encodes $N$ spatial nodes across $T$ time steps into a $d$-dimensional vector space. Storing or retrieving such embeddings indexing $O(N\cdot T\cdot d)$-sized tensors, where both spatial and temporal dependencies are entangled. This representation incurs large computational, memory, and I/O costs during retrieval. Inspired by low-rank approximation techniques such as LoRA (hu2021loralowrankadaptationlarge; ruan2025stloralowrankadaptationspatiotemporal), we instead decouple the embedding into two compact components: a temporal representation $U\in\mathbb{R}^{T\times d}$ and a spatial representation $V\in\mathbb{R}^{N\times d}$ such that:
$H\approx UV^{\top},\quad U\in\mathbb{R}^{T\times d},\;V\in\mathbb{R}^{N\times d},$
This decomposition reduces the storage and retrieval complexity from $O(N\cdot T\cdot d)$ down to $O((N+T)\cdot r)$, where $r$ is the disentangled feature rank. Therefore, separating time and space directly lowers computational complexity and improves retrieval efficiency, while maintaining expressivity through low-rank factorization.
Dual-dimensional retrieval also simplifies the maintenance of memory stores. Instead of building a monolithic database of high-dimensional spatio-temporal vectors, we maintain two retrieval stores: one specialized for temporal embeddings $\{U\}$, and one for spatial embeddings $\{V\}$. This allows dimension-specific indexing (e.g., FAISS similarity search) to scale more efficiently and robustly with dataset size, as each store grows linearly with $k\cdot T$ or $k\cdot N$ rather than with $k^{2}\cdot N\times T$ as $k$ is the selection coefficient. Moreover, by analogy to bilinear or Kronecker product structures, the fusion of retrieved $U$ and $V$ embeddings reconstructs the original spatio-temporal interactions, thereby preserving coupled dependencies. In summary, the dual-dimensional design not only reduces overhead and simplifies database operations, but also retains the ability to capture high-dimensional spatio-temporal structures in a manner consistent with established low-rank modeling approaches.
Appendix G Limitation and Future Work
@@
Appendix H Social Impact
Our proposed method has the potential to generate significant societal value by advancing the capabilities of Intelligent Transportation Systems (ITS). Efficient and accurate traffic forecasting enables applications such as congestion mitigation, dynamic route optimization, and demand-aware public transportation scheduling. By explicitly augmenting spatio-temporal models with retrieval mechanisms, RAST enhances fine-grained prediction accuracy under complex and heterogeneous traffic dynamics while maintaining computational efficiency. This ability enables city-level transportation agencies to make proactive, data-driven decisions that enhance mobility, reduce carbon emissions associated with traffic congestion, and facilitate timely responses to unexpected incidents, thereby promoting sustainability and safety in urban environments.
Beyond ITS, our framework can be generalized to other spatio-temporal domains where long-range dependencies and heterogeneity are critical, including climate forecasting, energy demand modeling, and healthcare resource allocation. In these domains, improved forecasting accuracy directly contributes to social welfare by enhancing preparedness, resource efficiency, and resilience against uncertainty. In conclusion, RAST not only addresses methodological challenges in traffic prediction but also provides a universal paradigm that can be adapted to broader spatio-temporal applications with tangible social impact.